# ClinNoteAgents: An LLM Multi-Agent System for Predicting and Interpreting Heart Failure 30-Day Readmission from Clinical Notes

### v1

@@ ABSTRACT
Heart failure (HF) is one of the leading causes of rehospitalization among older adults in the United States. Although clinical notes contain rich, detailed patient information and make up a large portion of electronic health records (EHRs), they remain underutilized for HF readmission risk analysis. Traditional computational models for HF readmission often rely on expert-crafted rules, medical thesauri, and ontologies to interpret clinical notes, which are typically written under time pressure and may contain misspellings, abbreviations, and domain-specific jargon. We present ClinNoteAgents, an LLM-based multi-agent framework that transforms free-text clinical notes into (1) structured representations of clinical and social risk factors for association analysis and (2) clinician-style abstractions for HF 30-day readmission prediction. We evaluate ClinNoteAgents on 3,544 notes from 2,065 patients (readmission rate=35.16%), demonstrating strong performance in extracting risk factors from free-text, identifying key contributing factors, and predicting readmission risk. By reducing reliance on structured fields and minimizing manual annotation and model training, ClinNoteAgents provides a scalable and interpretable approach to note-based HF readmission risk modeling in data-limited healthcare systems.
@@ INTRODUCTION
Heart failure (HF) remains a major global health challenge, affecting more than 55 million individuals worldwide, with nearly 80% of cardiovascular deaths occurring in low- and middle-income nations yan2021GBDHF. Approximately 25% of HF patients are readmitted within 30 days khan2021HFReadmission, imposing substantial clinical and financial burdens on health systems. Multiple studies have identified diverse contributors to 30-day readmission of HF, including HF exacerbationEltelbany2019CausesHFReadmission, comorbidities such as chronic obstructive pulmonary disease, chronic kidney disease, anemia infection Jain2023EtiologiesPredictors30DayHFReadmission, and socioeconomic factorsReddy2019ReadmissionsHF. Given the complex, multi-factorial nature of HF, computational risk modeling for HF readmission often requires comprehensive longitudinal patient data butler2012hospital; umehara2020factors. Electronic health record (EHR) data, which capture demographics, diagnoses, laboratory results, and medications, have therefore become a central data source for studying HF outcomes, including 30-day readmission mahajan2017validated; yu2024machine. While containing valuable multimodal health information for HF patients, EHRs are often incomplete or unavailable in developing countries murray2022GBD due to financial, technological, and organizational barriers bostan2024OpenSourceEHR. This challenge is particularly evident in developing countries in Asia and Africa. Many hospitals in Bangladesh and Indonesia continue to rely on handwritten or locally stored digital notes due to financial constraints, limited IT capacity, and poor interoperability infrastructure Hossain2025EMR; Taher2025HRIS. Similarly, healthcare facilities in Kenya, Uganda, and Ghana frequently experience unstable internet connectivity and shortages of trained health informatics personnel Alzghaibi2025EHRS; Mensah2024EHR. In these settings, unstructured clinical notes often serve as the primary source of documented patient information. Despite widespread adoption of EHR in the U.S., around 80% of clinical information remains embedded in free-text notes Kong2019ManagingUnstructured. Therefore, clinical notes offer a pragmatic and scalable strategy for constructing predictive models when access to structured EHR data is limited.
@@
Early HF readmission models relied primarily on structured EHR data such as demographics, socioeconomic status, medical history, and laboratory measurements Frizzell2017; Pishgar2022. More recent work has incorporated unstructured clinical notes to capture richer contextual and temporal information Liu2019readmission, with hybrid models combining structured variables and note-derived embeddings yielding further gains Golas2018. This shift underscores the potential of text-based modeling for early detection of readmission risk. Growing evidence also highlights the importance of social determinants of health (SDOHs), defined by the World Health Organization as the conditions in which individuals live and work whoSDOH2025. SDOHs have been repeatedly linked to HF outcomes, with employment status, housing stability, and social support identified as major contributors to readmission risk.guevara2024llmsdoh. However, SDOHs are rarely structured in EHRs, limiting their use in predictive systems. Recent advances in natural language processing (NLP) and large language models (LLMs) have enabled the extraction of both clinical and social risk factors from unstructured notes, with machine learning and transformer models achieving state-of-the-art performance in HF readmission prediction using discharge notes alnomasy2025HFReadmission; huang2020ClinicalBERT. Summarization-based preprocessing is shown to further enhance predictive signal quality boll2025DistillNote, and recent studies show that LLMs can extract SDOH with near-clinician accuracy consoli2025SDoHGPT; gu2025SBDHReader. Despite these advances, most existing approaches rely on predefined SDOH taxonomies or focus solely on social factors, often overlooking clinical predictors for readmission riskshao2025MiningSDOH. These gaps highlight the need for a unified framework that jointly extracts and harmonizes social and clinical determinants from discharge notes to enable scalable readmission modeling.
@@
To address these limitations, we present ClinNoteAgents, a LLM-based modular framework that transforms unstructured discharge notes into structured, interpretable representations of clinical and social risk factors for HF 30-day readmission. Our framework integrates (1) a structural extractor of clinical and social risk factors for statistical analysis of their relationship to readmission outcomes, and (2) a summarizer that produces concise qualitative or mixed-evidence summaries for readmission predictive modeling. By reducing reliance on structured EHR and minimizing manual annotation, ClinNoteAgents enables scalable HF risk analysis from discharge documentation in data-limited healthcare settings.
@@
This study used the publicly available Medical Information Mart for Intensive Care III (MIMIC-III) database johnson2016mimic. Following established conventions in prior studies shao2025MiningSDOH, heart failure (HF) patients were identified using ICD-9 diagnosis codes (398.91, 402.01, 402.11, 402.91, 404.01, 404.03, 404.11, 404.13, 404.91, 404.93, and all codes beginning with 428). For each patient with multiple hospitalizations, we constructed readmission pairs by linking each index admission with its subsequent readmission. The discharge summary from the earlier admission served as the input note $\mathbf{X}_{i}$ for modeling, while the outcome label $y_{i}$ was determined by the time interval between discharge and the next admission ($y_{i}=1$ if the interval $\leq 30$ days; otherwise $y_{i}=0$). Table 1 contains the summary statistics of the HF cohort used in this study.
@@
We propose ClinNoteAgents, an LLM-based multi-agent system for comprehensive clinical notes analytics. Our system comprises three agents: the Risk Factor Extractor, the Risk Factor Normalizer, and the Note Summarizer. The overall model framework is outlined in Figure 1. All LLM agents were implemented using Qwen3-14B yang2025qwen3 with thinking mode enabled. The extractor extracts clinical and SDOH variables; the normalizer and labeler standardize heterogeneous SDOH expressions into LLM-derived categories; and the summarizer produces clinician-style abstracts for downstream modeling.
@@
Prompt example for LLM structural extractor You are a clinical expert in extracting structured information from discharge summaries. Extract only the specific clinical factors that are important for predicting 30-day readmission in heart failure (HF) patients. Output Schema Example: { {internallinenumbers*} "Charted_SDOHs": {"Gender": "M", "Age": "50", "Language": "null", "Marital_Status": "Married"}, "NonCharted_SDOHs": { {internallinenumbers*} "Alcohol_use": "No","Tobacco_use": "1 ppd x 35y, quit 3 months ago","Drug_use": "No","Transportation": "null", "Housing": "null", "Parental": "null", "Employment_Status": "Retired", "Social_Support": "null" }, "Clinical_Info": { "Vitals": { {internallinenumbers*} "Body_Temperature": "96 F", "Heart_Rate": "90 bpm", "Respiration_Rate": "18", "Blood_Pressure": "120/60", "SpO2": "99%", "Weight": "100 kg", "Height": "180 cm" } }, "Chief_Complaint": { "Symptoms": "...", "Description": "..." }, "Diagnoses": [ {"Condition": "Renal cell cancer", "Details": ""}, ... ] }
Risk factor normalizer. Both charted and uncharted SDOH were often documented in heterogeneous free-text, limiting their use in quantitative analysis. To address this, we developed an LLM-based normalization module that standardized these entries into categorical values through a two-stage process: the first LLM (the normalizer) generated a concise set of normalized categories for each variable, and the second LLM (the labeler) assigned each extracted value to one of these categories. For charted SDOHs, only language and marital status required normalization, as gender was already standardized (male and female). All uncharted SDOHs underwent the same normalization pipeline. Results were compared to existing reference taxonomies (MIMIC-SBDH ahsan2021mimicsbdh and LLM-SDOH guevara2024llmsdoh) using predefined categories.
@@
Prompt example (no-number summary) You are a clinical expert… Given the discharge summary, produce a summary of the most noticeable abnormal findings, symptoms, and risk indicators that could be related to 30-day hospital readmission for heart failure (HF) patients. Transform the discharge summary into a clinician-style qualitative abstraction, preserving section headers, while removing numbers and converting them to qualitative descriptors, focused on the potential 30-day readmission risk for HF patients.
@@ RESULTS
We evaluated the risk-factor mining pipeline across its three components: (1) extraction, (2) normalization, and (3) correlation analysis. Risk-factor mining entails transforming each discharge note into structured clinical and social variables and assessing their correlations with HF 30-day readmission. First, we assessed the extraction fidelity for clinical variables and charted SDOH using structured EHR as a surrogate ground truth; and we used an LLM-as-a-judge framework huggingface2024llmjudge for diagnoses. Next, we evaluated the LLM-based normalization module, which standardizes heterogeneous SDOH expressions into analyzable categories. Finally, we conducted statistical association analyses to identify risk factors significantly correlated with HF readmission.
Structured EHR as surrogate ground truth. Clinical measurements (vitals) and charted SDOH were evaluated against corresponding EHR tables to assess extraction coverage and conditional accuracy. Coverage, defined as the proportion of patients with non-null LLM-extracted values (“% Extracted” in Table 2), ranged from $4.03\%$ for Height to $89.25\%$ for Heart Rate. Among SDOHs, Gender achieved the highest coverage ($99.18\%$), whereas Language and Marital Status were less frequently detected ($6.07\%$ and $25.73\%$, respectively). For each non-null extraction, values were compared with ground truth within variable-specific tolerance ranges (“Tolerance Range” in Table 2) to compute conditional accuracy (“Cond Acc” in Table 2). Mixed-unit variables were compared in their native units when aligned. Otherwise, both were converted to canonical units ($^{\circ}\mathrm{C}$, cm, kg) before comparison. For each vital within an admission, extracted and ground-truth values were converted to canonical units, aggregated, and summarized using medians. Mean absolute error and mean absolute percentage error (“MAE” and “MAPE” in Table 2) were computed without tolerance adjustments to reflect raw deviation magnitudes.
Table 2: Agreement with structured EHR ground truth for extracted clinical variables and charted SDOH
@@
Evaluation via LLM-as-a-judge. Diagnosis extraction was evaluated using an LLM-as-a-judge framework huggingface2024llmjudge, comparing LLM-extracted diagnoses with associated ICD codes. For each patient, the judge assigned a score from 0 (lowest) to 5 (highest) based on the semantic similarity between the extracted diagnoses and the ICD-9 codes. We computed two metrics: conditional accuracy, the proportion of correctly identified diagnoses among those extracted by the LLM, and absolute accuracy, the proportion of all ICD-9 diagnoses correctly recovered. As summarized in Table 3, the LLM extracted fewer diagnoses per patient than the structured ICD-9 list. The average similarity score was 3.04, and the conditional accuracy was 62.27%.
Table 3: LLM-as-a-judge evaluation of diagnosis extraction
@@
Table 4: LLM-normalized charted and uncharted SDOH categories
@@
Table 5: Logistic regression correlation analysis for LLM-extracted clinical risk factors
@@
Table 6: Chi-square test results for LLM-extracted charted and uncharted SDOH
@@ DISCUSSION
LLM Risk Factor Extractor. Accuracy within tolerance was high across most vitals. Conditional accuracy was highest for cardiorespiratory measures—SpO2 (94.16%), Respiratory Rate (93.77%), and Blood Pressure ($\geq$90%)—whereas anthropometric variables such as Height (68.50%) and Weight (57.90%) performed noticeably worse. These reductions were likely caused by heterogeneous reporting formats in discharge notes. Height may be recorded in centimeters or feet/inches, and weight in kilograms or pounds. Such variations may increase unit-conversion errors. Because LLMs exhibit unstable quantitative reasoning and units in free-text form cannot be reliably validated, LLM-driven unit conversion was minimized during extraction. Nevertheless, values in uncommon units may still be misinterpreted. These patterns are consistent with evidence that LLMs hallucinate unsupported or incorrect clinical details during information extraction from discharge notes. das2025hallucination. MAPE remained below 15% for most vitals, indicating close agreement with the EHR surrogate ground truth. Small deviations occurred for Temperature (1.40%) and SpO2 (2.77%), while more variable measures such as Blood Pressure (13.24%-19.99%) and Respiration Rate (19.14 %) showed larger discrepancies, consistent with their expected within-admission fluctuation. The LLM-as-a-judge evaluation showed that the model extracted fewer diagnoses per patient than the ICD-9 documentation. This is consistent with prior works that showed discrepancies between ICD coding and discharge notes kim2021icdaccuracy. Despite the difference, the LLM achieved a moderately strong similarity score (mean 3.04) and a conditional accuracy of 62.67%. This suggests that the extracted diagnoses were often semantically aligned with the ICD codes. These findings indicate LLM-based diagnosis extraction is feasible for scaling chart reviews, though discrepancies between clinical notes and coding systems warrant caution when using extracted diagnoses for downstream modeling.
@@
LLM Normalization and Correlation Analysis. The LLM-based normalization agent enabled correlation analysis that was otherwise infeasible on raw extractions. Prior works typically map SDOHs onto predefined taxonomies such as MIMIC-SBDH ahsan2021mimicsbdh and LLM-SDoH guevara2024llmsdoh. However, these schemes are coarse and often collapse clinically distinct expressions into a small set of categories (e.g., binary community factors, single “education” flag, and limited alcohol/tobacco categories). This limits cohort-specific nuance and may attenuate statistical associations. In contrast, the LLM-generated categories provided concise and cohort-specific representations while preserving relevant contextual detail. Logistic regression identified age, weight, and blood pressure (BP) as significantly associated with HF readmission. Age showed a positive association, consistent with established epidemiology in which older patients are at higher risk of readmission shao2025MiningSDOH. Systolic and diastolic BP exhibited negative associations, aligning with reports that low BP is associated with acute heart failure kim2024lowBP. Lower discharge weight may similarly reflect HF risk. The observed correlations underscore the potential of LLM-based extraction to support early identification and monitoring of high-risk patients. Other vital signs were not significant, suggesting that their discharge-time values provide limited discriminative value for readmission. Chi-square analysis found housing to be the only significant SDOH. Other social factors commonly reported as influential, including smoking, social support, and marital statuscalvillo2013socialsupport, were not significant in our cohort. These discrepancies reflect the under-documentation of SDOH in discharge notes or limitations in LLM extraction, highlighting the challenge of using discharge notes alone to capture SDOH.
@@
30-Day Readmission Predication using LLM Summaries. LLM-generated summaries preserved most of the predictive signal needed for HF readmission modeling. Although summarization did not improve classification performance as we initially expected, performance declines were modest across all models, even with 60–90% text reduction. This suggests that the LLM preserved the majority of risk-relevant information, such as comorbidities, discharge stability, and key physiological measurements, while removing redundant content and noise. Substantial compression reduces token requirements for transformer models and produces more concise, interpretable inputs for downstream review. Our results demonstrate that LLM summarization offers a balance between information preservation and computational efficiency, enabling scalable representation of clinical text while maintaining clinically meaningful signals for risk prediction.
@@
Limitations. One limitation of our study is the indirect evaluation of LLM-extracted risk factors. For SDOH variables and most vital signs, we relied on corresponding structured EHR fields as surrogate ground truth and assessed extraction quality using tolerance-based accuracy and an LLM-as-judge approach. Another limitation is that we did not observe substantial performance gains from summarization alone. Finally, evaluating patients across multiple sites may be needed to demonstrate the consistency and generalizability of the proposed multi-agent system.

### v2

@@ ABSTRACT
Heart failure (HF) is one of the leading causes of rehospitalization among older adults in the United States. Although clinical notes contain rich, detailed patient information and make up a large portion of electronic health records (EHRs), they remain underutilized for HF readmission risk analysis. Traditional computational models for HF readmission often rely on expert-crafted rules, medical thesauri, and ontologies to interpret clinical notes, which are typically written under time pressure and may contain misspellings, abbreviations, and domain-specific jargon. We present ClinNoteAgents, an LLM-based multi-agent framework that transforms free-text clinical notes into (1) structured representations of clinical and social risk factors for association analysis and (2) clinician-style abstractions for HF 30-day readmission prediction. We evaluate ClinNoteAgents on 3,544 notes from 2,065 patients (readmission rate=35.16%), demonstrating high extraction fidelity for clinical variables (conditional accuracy $\geq$90% for multiple vitals), key risk factor identification, and preservation of predictive signal despite 60–90% text reduction. By reducing reliance on structured fields and minimizing manual annotation and model training, ClinNoteAgents provides a scalable and interpretable approach to note-based HF readmission risk modeling in data-limited healthcare systems.
@@ INTRODUCTION
Heart failure (HF) remains a major global health challenge, affecting more than 55 million individuals worldwide, with nearly 80% of cardiovascular deaths occurring in low- and middle-income nations 37. Approximately 25% of HF patients are readmitted within 30 days 21, imposing substantial clinical and financial burdens on health systems. Multiple studies have identified diverse contributors to readmission of HF, including HF exacerbation11, comorbidities such as chronic obstructive pulmonary disease, chronic kidney disease 19, and socioeconomic factors30. Given the complex, multi-factorial nature of HF, computational risk modeling for HF readmission often requires comprehensive longitudinal patient data 34. Electronic health record (EHR) data, which capture demographics, diagnoses, laboratory results, and medications, have therefore become a central data source for studying HF outcomes, including 30-day readmission 26; 38. While containing valuable multimodal health information for HF patients, EHRs are often incomplete or unavailable in developing countries 28 due to financial, technological, and organizational barriers 7. This challenge is particularly evident in developing countries in Asia and Africa. Many hospitals in Bangladesh and Indonesia continue to rely on handwritten or locally stored digital notes due to financial constraints, limited IT capacity, and poor interoperability infrastructure 17; 32. Similarly, healthcare facilities in Kenya, Uganda, and Ghana frequently experience unstable internet connectivity and shortages of trained health informatics personnel 5; 27. In these settings, unstructured clinical notes often serve as the primary source of documented patient information. Despite widespread adoption of EHR in the U.S., around 80% of clinical information remains embedded in free-text notes 24. Therefore, clinical notes offer a pragmatic strategy for constructing predictive models when access to structured EHR data is limited.
@@
Early HF readmission models relied primarily on structured EHR data such as demographics, socioeconomic status, medical history, and laboratory measurements 13; 29. More recent work has incorporated unstructured clinical notes to capture richer contextual and temporal information 25, with hybrid models combining structured variables and note-derived embeddings yielding further gains 14. This shift underscores the potential of text-based modeling for early detection of readmission risk. Growing evidence also highlights the importance of social determinants of health (SDOHs), defined by the World Health Organization as the conditions in which individuals live and work 36. SDOHs have been repeatedly linked to HF outcomes, with employment status, housing stability, and social support identified as major contributors to readmission risk.16. However, SDOHs are rarely structured in EHRs, limiting their use in predictive systems. Recent advances in natural language processing (NLP) and large language models (LLMs) have enabled the extraction of both clinical and social risk factors from unstructured notes, with machine learning and transformer models achieving state-of-the-art performance in HF readmission prediction using discharge notes 4; 18. Summarization-based preprocessing is shown to further enhance predictive signal quality 6, and recent studies show that LLMs can extract SDOH with near-clinician accuracy 9; 15. Despite these advances, most existing approaches rely on predefined SDOH taxonomies or focus solely on social factors, often overlooking clinical predictors for readmission risk31. These gaps highlight the need for a unified framework that jointly extracts and harmonizes social and clinical determinants from discharge notes to enable scalable readmission modeling.
@@
To address these limitations, we present ClinNoteAgents, a LLM-based modular framework that transforms unstructured discharge notes into structured, interpretable representations of clinical and social risk factors for HF 30-day readmission. Our framework integrates (1) a structural extractor of clinical and social risk factors for statistical analysis of their relationship to readmission outcomes, and (2) a summarizer that produces qualitative or mixed-evidence summaries for predictive modeling. By reducing reliance on structured EHR and minimizing manual annotation, ClinNoteAgents enables scalable HF risk analysis from discharge documentation in data-limited healthcare settings.
@@
This study used the publicly available Medical Information Mart for Intensive Care III (MIMIC-III) database 20. Following established conventions in prior studies 31, heart failure (HF) patients were identified using ICD-9 diagnosis codes (398.91, 402.01, 402.11, 402.91, 404.01, 404.03, 404.11, 404.13, 404.91, 404.93, and all codes beginning with 428). For each patient with multiple hospitalizations, we constructed readmission pairs by linking each index admission with its subsequent readmission. The discharge summary from the earlier admission served as the input note $\mathbf{X}_{i}$ for modeling, while the outcome label $y_{i}$ was determined by the time interval between discharge and the next admission ($y_{i}=1$ if the interval $\leq 30$ days; otherwise $y_{i}=0$). Table 1 summarizesthe cohort used in this study.
@@
We propose ClinNoteAgents, an LLM-based multi-agent system for comprehensive clinical notes analytics. Our system comprises three agents: the Risk Factor Extractor, the Risk Factor Normalizer, and the Note Summarizer. The overall model framework is outlined in Figure 1. All LLM agents were implemented using Qwen3-14B 33 with thinking mode enabled. The extractor extracts clinical and SDOH variables; the normalizer and labeler standardize heterogeneous SDOH expressions into LLM-derived categories; and the summarizer produces clinician-style abstracts for downstream modeling.
@@
Prompt example for LLM structural extractor You are a clinical expert in extracting structured information from discharge summaries. Extract only the specific clinical factors that are important for predicting 30-day readmission in heart failure (HF) patients. Output Schema Example: { {internallinenumbers*} "Charted_SDOHs": {"Gender": "M", "Age": "50", "Language": "null", "Marital_Status": "Married"}, {internallinenumbers*} "NonCharted_SDOHs": {"Alcohol_use": "No","Tobacco_use": "1 ppd x 35y, quit 3 months ago","Drug_use": "No","Transportation": "null", "Housing": "null", "Parental": "null", "Employment_Status": "Retired", "Social_Support": "null"}, "Clinical_Info": { {internallinenumbers*} "Vitals": {"Body_Temperature": "96 F", "Heart_Rate": "90 bpm", "Respiration_Rate": "18", "Blood_Pressure": "120/60", "SpO2": "99%", "Weight": "100 kg", "Height": "180 cm"}}, "Chief_Complaint": {"Symptoms": "...", "Description": "..."}, "Diagnoses": [{"Condition": "Renal cell cancer", "Details": "..."}, ...] }
Risk factor normalizer. Both charted and uncharted SDOH were often documented in heterogeneous free-text, limiting their use in quantitative analysis. To address this, we developed an LLM-based normalization module that standardized these entries into categorical values through a two-stage process: the first LLM (the normalizer) generated a concise set of normalized categories for each variable, and the second LLM (the labeler) assigned each extracted value to one of these categories. For charted SDOHs, only language and marital status required normalization, as gender was already standardized (male and female). All uncharted SDOHs underwent the same normalization pipeline. Results were compared to existing reference taxonomies (MIMIC-SBDH 3 and LLM-SDOH 16) using predefined categories.
@@
Prompt example (no-number summary) [Same as overall summary above]. Transform the discharge summary into a clinician-style qualitative abstraction, preserving section headers, while removing numbers and converting them to qualitative descriptors, focused on the potential 30-day readmission risk for HF patients.
@@ RESULTS
We evaluated the risk-factor mining pipeline across extraction, normalization, and correlation analysis. Risk-factor mining transforms each discharge note into structured clinical and social variables and assesses their correlations with HF readmission. We first assessed extraction fidelity for clinical variables and charted SDOH using structured EHR as a surrogate ground truth, and applied an LLM-as-a-judge framework 12 for diagnoses. Next, we evaluated the LLM-based normalization module, which standardizes heterogeneous SDOH expressions into analyzable categories. Finally, we conducted statistical association analyses to identify risk factors significantly correlated with HF readmission.
Structured EHR as surrogate ground truth. Clinical measurements (vitals) and charted SDOH were evaluated against corresponding EHR tables to assess extraction coverage and conditional accuracy. Coverage, defined as the proportion of patients with non-null LLM-extracted values (“% Extracted” in Table 2), ranged from $4.03\%$ for Height to $89.25\%$ for Heart Rate. Among SDOHs, Gender achieved the highest coverage ($99.18\%$), whereas Language and Marital Status were less frequently detected ($6.07\%$ and $25.73\%$, respectively). For each non-null extraction, values were compared with ground truth within variable-specific tolerance ranges (“Tolerance Range” in Table 2) to compute conditional accuracy (“Cond Acc” in Table 2). Tolerance ranges were designed to account for variability in clinical documentation, including rounding differences, unit conversions, or discrepancies between EHR and discharge notes. Mixed-unit variables, namely Temperature, Height, and Weight, were compared in their native units when aligned. When discrepancies occurred, both LLM-extracted and structured EHR values were converted to canonical units ($^{\circ}\mathrm{C}$, cm, and kg) before comparison. Other vitals were consistently recorded and did not require unit harmonization. For each vital within an admission, extracted and ground-truth values were converted to canonical units, aggregated, and summarized using medians. Mean absolute error and mean absolute percentage error (“MAE” and “MAPE” in Table 2) were computed without tolerance adjustments to reflect raw deviation magnitudes.
Table 2: Agreement with structured EHR ground truth for extracted clinical variables and charted SDOH.
@@
Evaluation via LLM-as-a-judge. Diagnosis extraction was evaluated using an LLM-as-a-judge framework 12, comparing LLM-extracted diagnoses with associated ICD codes. For each patient, the judge assigned a score from 0 (lowest) to 5 (highest) based on the semantic similarity between the extracted diagnoses and the ICD-9 codes. We computed two metrics: conditional accuracy, the proportion of correctly identified diagnoses among those extracted by the LLM, and absolute accuracy, the proportion of all ICD-9 diagnoses correctly recovered. As summarized in Table 3, the LLM extracted fewer diagnoses per patient than the structured ICD-9 list. The average similarity score was 3.04, and the conditional accuracy was 62.27%.
Table 3: LLM-as-a-judge evaluation of diagnosis extraction.
@@
Table 4: LLM-normalized charted and uncharted SDOH categories.
@@
Table 5: Logistic regression correlation analysis for LLM-extracted clinical risk factors.
@@
Table 6: Chi-square test results for LLM-extracted charted and uncharted SDOH.
@@ DISCUSSION
LLM Risk Factor Extractor. Accuracy within tolerance was high across most vitals. Conditional accuracy was highest for cardiorespiratory measures—SpO2 (94.16%), Respiratory Rate (93.77%), and Blood Pressure ($\geq$90%)—whereas anthropometric variables such as Height (68.50%) and Weight (57.90%) performed noticeably worse. These reductions were likely caused by heterogeneous reporting formats in discharge notes. Height may be recorded in centimeters or feet/inches, and weight in kilograms or pounds. Such variations may increase unit-conversion errors. Because LLMs exhibit unstable quantitative reasoning, and units in free-text form cannot be reliably validated, LLM-driven unit conversion was minimized during extraction. Nevertheless, values in uncommon units may still be misinterpreted. These patterns are consistent with evidence that LLMs hallucinate unsupported or incorrect clinical details during information extraction from discharge notes. 10. MAPE remained below 15% for most vitals, indicating close agreement with the EHR surrogate ground truth. Small deviations occurred for Temperature (1.40%) and SpO2 (2.77%), while more variable measures such as Blood Pressure (13.24%-19.99%) and Respiration Rate (19.14 %) showed larger discrepancies, consistent with their expected within-admission fluctuation. Several vitals and SDOHs showed low extraction coverage, largely due to under-documentation in discharge notes rather than extraction failure. Association analyses were therefore conducted only on patients with non-null values for each variable. While this approach may introduce selection effects, the findings should be interpreted as exploratory signals rather than population-level estimates. The LLM-as-a-judge evaluation showed that the model extracted fewer diagnoses per patient than the ICD-9 documentation. This is consistent with prior works that showed discrepancies between ICD coding and discharge notes 22. Despite the difference, the LLM achieved a moderately strong similarity score (mean 3.04) and a conditional accuracy of 62.67%. This suggests that the extracted diagnoses were often semantically aligned with the ICD codes. These findings indicate LLM-based diagnosis extraction is feasible for scaling chart reviews, though discrepancies between clinical notes and coding systems warrant caution when using extracted diagnoses for downstream modeling.
@@
LLM Normalization and Correlation Analysis. The LLM-based normalization agent enabled correlation analysis that was otherwise infeasible on raw extractions. Prior works typically map SDOHs onto predefined taxonomies such as MIMIC-SBDH 3 and LLM-SDoH 16. However, these schemes are coarse and often collapse clinically distinct expressions into a small set of categories (e.g., binary community factors and limited alcohol/tobacco categories). This limits cohort-specific nuance and may attenuate statistical associations. In contrast, the LLM-generated categories provided concise and cohort-specific representations while preserving relevant contextual detail. Logistic regression identified age, weight, and blood pressure (BP) as significantly associated with HF readmission. Age showed a positive association, consistent with established epidemiology in which older patients are at higher risk of readmission 31. Systolic and diastolic BP exhibited negative associations, aligning with reports that low BP is associated with acute heart failure 23. Lower discharge weight may similarly reflect HF risk. The observed correlations underscore the potential of LLM-based extraction to support early identification and monitoring of high-risk patients. Other vital signs were not significant, suggesting that their discharge-time values provide limited discriminative value for readmission. Chi-square analysis found housing to be the only significant SDOH. Other social factors commonly reported as influential, including smoking, social support, and marital status8, were not significant in our cohort. These discrepancies reflect the under-documentation of SDOH in discharge notes or limitations in LLM extraction, highlighting the challenge of using discharge notes alone to capture SDOH.
@@
30-Day Readmission Predication using LLM Summaries. LLM-generated summaries preserved most predictive signal needed for HF readmission modeling. Although summarization did not improve classification performance as initially expected, performance declines were modest across models, even with 60–90% text reduction. This suggests the LLM preserved the majority of risk-relevant information, such as comorbidities, discharge stability, and key physiological measurements, while removing redundant content and noise. Substantial compression reduces token requirements for transformer models and produces more concise inputs for downstream review. Our results demonstrate that LLM summarization offers a balance between information preservation and computational efficiency, enabling scalable representation of clinical text while maintaining clinically meaningful signals for risk prediction.
Model Choice in ClinNoteAgents. We selected LLM agents over pre-trained language models (PLMs) such as ClinicalBERT because our note-to-structure tasks (extraction, normalization, and summarization) require generative reasoning and minimal supervision. LLMs are known to extract clinical entities and numeric data from unstructured clinical notes with high fidelity in zero- and few-shot settings 2; 1, enabling direct extraction of heterogeneous risk factors from discharge notes. LLMs also support SDOH extraction 16 and, unlike PLMs, can perform normalization without requiring task-specific pretraining 15. Moreover, LLMs provide substantially longer context windows than PLMs and is able to achieve summarization quality approaching expert benchmarks. 35. We employed Qwen3 across agents to balance computing efficiency and performance, given its strong results in clinical information extraction benchmarks. 10.
@@
Limitations. One limitation of our study is the indirect evaluation of LLM-extracted risk factors. For SDOH variables and most vital signs, we relied on structured EHR fields as surrogate ground truth and assessed extraction quality using tolerance-based accuracy and an LLM-as-judge approach. Another limitation is that we did not observe substantial performance gains from summarization alone. Finally, we did not include clinician-led evaluations of LLM extraction or summarization results, which limits formal assessment of clinical fidelity and real-world applicability.